# ERDŐS–PÓSA PROPERTY OF CYCLES THAT ARE FAR APART

### v2

@@ Acknowledgements
Recall that $F_{(z)}$ is a rooted forest. Now for each $v\in V(F_{1})$ with $v>v_{z-1}$, we let
$\mathcal{B}_{z}(v)$ =$=\{A\in\mathcal{A}:\pi_{1}(A)\subseteq F_{(z)}(v)\},$ (12)
$\mathcal{C}_{z}(v)$ =$=\{(\pi_{2}(A),\ldots,\pi_{c}(A)):A\in\mathcal{B}_{z}(v)\}.$
If there exists $v\in V(F_{1})$ such that $v>v_{z-1}$ and either $c=1$, or $c\geq 2$ and $\mathcal{C}_{z}(v)$ contains an independent set of size $k^{c-1}$ then we define $v_{z}$ to be the minimum such $v$ (with respect to the ordering $\sigma$). If $c\geq 2$, we define the set $\mathcal{D}_{z}$ to be an independent set of size $k^{c-1}$ contained in $\mathcal{C}_{z}(v_{z})$. If no such vertex $v$ exists, then we set $m=z-1$ and stop the process.
We consider two separate cases: (1) $m\geq k$, in which case we show that $\mathcal{A}$ contains an independent set of size $k$, so (a) holds; (2) $m<k$, in which case we show that there exist $X_{1},\ldots,X_{c}$ that satisfy (b).
Suppose first that $m\geq k$ so that $v_{1},\ldots,v_{k}$ is well-defined. For each $z\in[k]$, define $\mathcal{B}_{z}:=\mathcal{B}_{z}(v_{z})$. Consider $\alpha,\beta\in[k]$ with $\alpha<\beta$. Since $v_{\alpha}<v_{\beta}$, we conclude that $v_{\beta}\not\in V(F_{1}(v_{\alpha}))$. In particular, $v_{\beta}\not\in V(F_{(\alpha)}(v_{\alpha}))$. Since $\alpha<\beta$ we have $v_{\alpha}\not\in V(F_{(\beta)})$. We conclude that
@@
By (12) for all $A\in\mathcal{B}_{\alpha}$, $B\in\mathcal{B}_{\beta}$, we have that $\pi_{1}(A)\subseteq F_{(\alpha)}(v_{\alpha})$, $\pi_{1}(B)\subseteq F_{(\beta)}(v_{\beta})$. Therefore by (13),
@@
If $c=1$, then taking one element in $\mathcal{B}_{z}$ for each $z\in[k]$ yields an independent set of size $k$, which establishes a. Now suppose $c\geq 2$ so, for each $z\in[k]$, we have that $\mathcal{D}_{z}$ is an independent family of $(F_{i})_{i\in\{2,\ldots,c\}}$-tuples and that $|\mathcal{D}_{z}|\geq k^{c-1}$. By 13, for each $z\in[k]$ there exists $D_{z}\in\mathcal{D}_{z}$ such that $\{D_{1},\ldots,D_{k}\}$ is independent.
@@
For each $z\in[k]$, select some $A_{z}\in\mathcal{B}_{z}$ such that $(\pi_{2}(A_{z}),\ldots,\pi_{c}(A_{z}))=D_{z}$. We now show that $\{A_{1},\ldots,A_{k}\}$ is an independent set of $(F_{i})_{i\in[c]}$-tuples. To this end, let $A:=A_{\alpha}$ and $B:=A_{\beta}$ for some $\alpha,\beta\in[k]$ with $\alpha<\beta$. We must show that $V(\pi_{i}(A))\cap V(\pi_{i}(B))=\emptyset$, for each $i\in[c]$. Since $A\in\mathcal{B}_{\alpha}$ and $B\in\mathcal{B}_{\beta}$, by (14) we have that $V(\pi_{1}(A))\cap V(\pi_{1}(B))=\emptyset$. Since $(\pi_{2}(A),\ldots,\pi_{c}(A))=D_{\alpha}$ and $(\pi_{2}(B),\ldots,\pi_{2}(B))=D_{\beta}$, and since $D_{\alpha}$ and $D_{\beta}$ are independent, we conclude that $V(\pi_{i}(A))\cap V(\pi_{i}(B))=\emptyset$, for each $i\in\{2,\ldots,c\}$. This proves that $\{A_{1},\ldots,A_{k}\}\subseteq\mathcal{A}$ is an independent set, so (a) holds.
@@
Now we consider the case in which $m<k$. If $c=1$ then taking $X_{1}:=\{v_{1},\ldots,v_{m}\}$ satisfies b with $\ell(k,1)=m\leq k-1$ and we are done. Now assume $c\geq 2$. For each $z\in[m]$, let $U_{z}$ be the set of (at most two) children of $v_{z}$ in $F_{(z)}$ and define $\mathcal{E}_{z}:=\bigcup_{u\in U_{z}}\mathcal{C}_{z}(u)$. By the minimality of $v_{z}$, $\mathcal{C}_{z}(u)$ does not contain an independent set of size $k^{c-1}$ for each $z\in[m]$ and $u\in U_{z}$. Therefore $\mathcal{E}_{z}$ does not contain an independent set of size $2k^{c-1}$ for each $z\in[m]$. Next, recall that $F_{1}$ is a tree with root $r_{1}$. If $r_{1}=v_{m}$, then define $\mathcal{E}_{m+1}:=\emptyset$. Otherwise, define $\mathcal{E}_{m+1}:=\mathcal{C}_{m+1}(r_{1})$. In the latter case, since the process for constructing $v_{1},\ldots,v_{m}$ stopped at iteration $m$, $\mathcal{E}_{m+1}=\mathcal{C}_{m+1}(r_{1})$ does not contain an independent set of size $k^{c-1}$. Therefore $\mathcal{A^{\prime}}:=\bigcup_{z\in[m+1]}\mathcal{E}_{z}$ does not contain an independent set of size $(m+1)2k^{c-1}\leq 2k^{c}$.
@@
Observe that $\mathcal{A}^{\prime}$ is a collection of non-trivial $(F_{i})_{i\in\{2,\ldots,c\}}$-tuples. Apply the inductive hypothesis with $c^{\prime}:=c-1$ and $k^{\prime}:=2k^{c}$ to the collection $\mathcal{A}^{\prime}$ of non-trivial $(F_{i})_{i\in\{2,\ldots,c\}}$-tuples. Since $\mathcal{A}^{\prime}$ does not contain an independent set of size $k^{\prime}$, this results in outcome b. That is, we obtain a sequence of sets $X_{2},\ldots,X_{c}$ with $X_{i}\subseteq V(F_{i})$ for each $i\in\{2,\ldots,c\}$ such that $\sum_{i=2}^{c}|X_{i}|\leq\ell(2k^{c},c-1)$, and, for each $(A_{2},\ldots,A_{c})\in\mathcal{A}^{\prime}$, there exists some $i\in\{2,\ldots,c\}$ such that $V(A_{i})\cap X_{i}\neq\emptyset$.
Let $X_{1}:=\{v_{1},\ldots,v_{m}\}$. We claim that $X_{1},\ldots,X_{c}$ satisfy the conditions of outcome b.
@@
First we show that, for each $(A_{1},\ldots,A_{c})\in\mathcal{A}$, there exists $i\in[c]$ such that $V(A_{i})\cap X_{i}\neq\emptyset$. If $(A_{2},\ldots,A_{c})\in\mathcal{A}^{\prime}$ then, by the inductive hypothesis, $V(A_{i})\cap X_{i}\neq\emptyset$ for some $i\in\{2,\ldots,c\}$. If $(A_{2},\ldots,A_{c})\notin\mathcal{A}^{\prime}$, then this is because there is no $z\in[m+1]$ such that $(A_{2},\ldots,A_{c})\in\mathcal{E}_{z}$. Let $v$ be the root of $A_{1}$. If $v$ has no $F_{1}$-ancestor in $X_{1}$, then $A_{1}\not\subseteq F_{(m+1)}(r_{1})$, which is the component of $F_{(m+1)}=F_{1}-X_{1}$ that contains $r_{1}$, because $(A_{2},\dots,A_{c})\notin\mathcal{E}_{m+1}=\mathcal{C}_{m+1}(r_{1})$. Since $v$ has no $F_{1}$-ancestor in $X_{1}$, $v$ is in $F_{(m+1)}(r_{1})$. Since $A_{1}\not\subseteq F_{(m+1)}(r_{1})$, it must be that $A_{1}$ contains a vertex in $X_{1}$, so $V(A_{1})\cap X_{1}\neq\emptyset$, as required. Otherwise, let $v_{z}$ be the ancestor of $v$ in $X_{1}$ that is closest to $v$. If $v=v_{z}$ then $V(A_{1})\cap X_{1}\supseteq\{v\}\supsetneq\emptyset$, as required. Otherwise $A_{1}$ is contained in $F_{1}(u)$ for some child $u$ of $v_{z}$. Since $A_{1}\notin\mathcal{E}_{z}\supseteq C_{z}(u)$, we have that $A_{1}\not\subseteq F_{(z)}(u)=F_{1}(u)-\{v_{1},\ldots,v_{z-1}\}$. Thus $V(A_{1})$ contains $v_{i}$ for some $i\in[z-1]$ so $V(A_{1})\cap X_{1}\supseteq\{v_{i}\}\supsetneq\emptyset$.
All that remains is to show that $\sum_{i\in[c]}|X_{i}|\leq\ell(k,c)$ for some $\ell:\mathbb{N}^{2}\to\mathbb{N}$. To do this, we see that
$\sum_{i\in[c]}|X_{i}|=|X_{1}|+\sum_{i=2}^{c}|X_{i}|=m+\sum_{i=2}^{c}|X_{i}|\leq m+\ell(2k^{c},c-1)<k+\ell(2k^{c},c-1)\enspace.$
Thus, $\ell(k,c)$ obeys the recurrence $\ell(k,c)\leq k+\ell(2k^{c},c-1)$ for $c\geq 1$ and $\ell(k,0)=0$. Clearly $\ell(k,c)$ depends only on $k$ and $c$, which is sufficient to establish the lemma. For a more precise estimate, we can show by induction on $c$ that $|X_{i}|\leq(\prod_{j=1}^{c-1}2^{j!})\,k^{c!}$ for each $i\in[c]$. By definition, $|X_{1}|=m\leq k$, which satisfies our requirement since $c\geq 1$. By the inductive hypothesis,
@@
for each $i\in\{2,\ldots,c\}$. Thus, $\sum_{i\in[c]}|X_{i}|\leq c\cdot(\prod_{j=1}^{c-1}2^{j!})\,k^{c!}$. ∎
Finally, we can complete the proof of Theorem 3 (restated here for convenience).
@@
By adding a vertex $r$ adjacent to one vertex in each component of $F$, we may assume that $F$ is a tree rooted at $r$. As in the proof of Lemma 15, we may further assume that $F$ is a binary tree.
Let $\ell:\mathbb{N}^{2}\to\mathbb{N}$ be the function given by Lemma 15.
The proof is by induction on $c$. When $c=1$, $\mathcal{A}$ is a collection of $(F_{i})_{i\in[1]}$-tuples and the result follows, with $\ell^{\star}(k,1)=\ell(k,1)$, from Lemma 15. Now assume that $c\geq 2$.
Let $A\in\mathcal{A}$ and let $c^{\prime}$ be the number of components of $A$. Thus $c^{\prime}\in[c]$. Define $\Pi(A):=(A_{1},\ldots,A_{c})$ so that $A_{1},\ldots,A_{c^{\prime}}$ are the components of $A$ while $A_{c^{\prime}+1},\ldots,A_{c}$ are the null graphs.
For each $\{i,j\}\in\binom{[c]}{2}$ and each $A\in\mathcal{A}$, define $A_{\{i,j\}}$ as the graph obtained from $A$ as follows. If one or both of $\pi_{i}(\Pi(A))$ or $\pi_{j}(\Pi(A))$ is the null graph then $A_{\{i,j\}}:=A$. Otherwise $A_{\{i,j\}}$ is the graph obtained from $A$ by adding the shortest path in $F$ that has one endpoint in $\pi_{i}(\Pi(A))$ and the other endpoint in $\pi_{j}(\Pi(A))$. In either case, $A_{\{i,j\}}$ has at most $c-1$ connected components.
For each $\{i,j\}\in\binom{[c]}{2}$, let $\mathcal{A}_{i,j}:=\{A_{\{i,j\}}:A\in\mathcal{A}\}$. Then $\mathcal{A}_{\{i,j\}}$ is a collection of non-null subgraphs of $F$ each having at most $c-1$ components. Now apply the inductive hypothesis to $\mathcal{A}_{\{i,j\}}$. If outcome a holds, then the $k$ pairwise vertex-disjoint members of $\mathcal{A}_{\{i,j\}}$ correspond to $k$ pairwise vertex-disjoint members of $\mathcal{A}$ and there is nothing more to do. Otherwise, outcome b holds and we obtain a set $X_{\{i,j\}}$ with $|X_{\{i,j\}}|\leq\ell^{\star}(k,c-1)$ such that $X_{\{i,j\}}\cap V(A_{\{i,j\}})\neq\emptyset$ for each $A\in\mathcal{A}$.
@@
For each $A\in\mathcal{A}_{0}$, define $\Pi_{0}(A)$ as the $(F_{i})_{i\in[m]}$-tuple in which $\pi_{i}(\Pi_{0}(A))$ is the component of $A$ contained in $F_{i}$ or the null graph if $V(A)\cap V(F_{i})=\emptyset$, for each $i\in[m]$. Since $A$ is non-null, $\Pi_{0}(A)$ is non-trivial. At this point, we can immediately apply Lemma 15 with $c^{\prime}:=m$ to the collection $\{\Pi_{0}(A):A\in\mathcal{A}_{0}\}$ of non-trivial $(F_{i})_{i\in[m]}$-tuples. If outcome a occurs then the resulting set of $k$ independent $(F_{i})_{i\in[m]}$-tuples corresponds to a set of pairwise vertex-disjoint elements of $\mathcal{A}_{0}\subseteq\mathcal{A}$, which satisfies the conditions of a. If outcome b occurs, then we obtain sets $X_{1},\ldots,X_{m}$ with $X_{i}\subseteq V(F_{i})$, for each $i\in[m]$, of total size at most $\ell(k,m)$ such that for each $A\in\mathcal{A}_{0}$, there exists $i\in[m]$ such that $V(A)\cap X_{i}\neq\emptyset$. Then $X:=X_{0}\cup\bigcup_{i\in[m]}X_{i}$ satisfies the conditions of b with
$|X|\leq|X_{0}|+|X_{1}|\leq\binom{c}{2}\,\ell^{\star}(k,c-1)+\ell(k,m)\leq\binom{c}{2}\,\ell^{\star}(k,c-1)+\ell(k,2\binom{c}{2}\,\ell^{\star}(k,c-1))\enspace.$
@@
which is sufficient to establish Theorem 3. However, with a bit more work, one can prove a bound of the form $\ell^{\star}(k,c)\in O(k^{h(c)})$ for some $h:\mathbb{N}\to\mathbb{N}$, which is polynomial in $k$ for any fixed $c$.
For each $I\in\binom{[m]}{c}$, define the set $\mathcal{A}_{I}:=\{A\in\mathcal{A}_{0}:A\subseteq\bigcup_{i\in I}F_{i}\}$. Since each element in $A$ has at most $c$ components and $V(A)\subseteq V(F_{0})$ for each $A\in\mathcal{A}_{0}$, each such $A$ appears in $\mathcal{A}_{I}$ for at least one $I\in\binom{[m]}{c}$. For each $I\in\binom{[m]}{c}$ and each $A\in\mathcal{A}_{I}$, define $\Pi_{I}(A):=(\pi_{i}(\Pi_{0}(A))_{i\in I}$ as the $(F_{i})_{i\in I}$-tuple obtained by restricting $\Pi_{0}(A)$ to the indices in $I$. For each $I\in\binom{[m]}{c}$, apply Lemma 15 to the collection $\{\Pi_{i}(A):A\in\mathcal{A}_{I}\}$ of $(F_{i})_{i\in I}$-tuples. If outcome a occurs for some $I\in\binom{[m]}{c}$ then the resulting set of $k$ independent $(F_{i})_{i\in I}$-tuples corresponds to a set of pairwise vertex-disjoint elements of $\mathcal{A}_{I}\subseteq\mathcal{A}$, which satisfies the conditions of a. If outcome b occurs for all $I\in\binom{[m]}{c}$, then we obtain a set $X_{I}\subseteq V(F_{0})\subseteq V(F)$ with $|X_{I}|\leq\ell(k,c)$ such that $V(A)\cap X_{I}\neq\emptyset$ for each $A\in\mathcal{A}_{I}$.

### v3

@@ 5 Proof of Theorem˜3
Recall that $F_{(z)}$ is a rooted forest and that, for each vertex $v$ of that forest, $F_{(z)}(v)$ denotes the subgraph of $F_{(z)}$ induced by all the descendants of $v$ in $F_{(z)}$ that is rooted at $v$. Now for each vertex $v$ of $F_{(z)}(r_{1})$ let
$\mathcal{B}_{z}(v)$ $=\{A\in\mathcal{A}:\pi_{1}(A)\subseteq F_{(z)}(v)\},$
$\mathcal{C}_{z}(v)$ $=\{(\pi_{2}(A),\ldots,\pi_{c}(A)):A\in\mathcal{B}_{z}(v)\}.$
If $c=1$ and $\mathcal{B}_{z}(r_{1})=\emptyset$, then we set m:=z1m:=z-1 and stop the process. If $c\geq 2$ and $\mathcal{C}_{z}(r_{1})$ contains no independent set of size $k^{c-1}$, then we set m:=z1m:=z-1 and stop the process. Otherwise, we define $v_{z}$ to be a minimal vertex in $F_{(z)}$ that satisfies one of the following properties:
  1. 1.
    $c=1$ and $\mathcal{B}_{z}(v)\neq\emptyset$;
  2. 2.
    $c\geq 2$ and $\mathcal{C}_{z}(v)$ contains an independent set of size $k^{c-1}$.
This completes the definition of the process.
We consider two separate cases: (1) $m\geq k$, in which case we show that $\mathcal{A}$ contains an independent set of size $k$, so (a) holds; (2) $m<k$, in which case we show that there exist $X_{1},\ldots,X_{c}$ that satisfy (a).
Suppose first that $m\geq k$ so that $v_{1},\ldots,v_{k}$ is well-defined. For each $z\in[k]$, define $\mathcal{B}_{z}:=\mathcal{B}_{z}(v_{z})$. Consider $\alpha,\beta\in[k]$ with $\alpha<\beta$. Note that $v_{\alpha}$ is not a vertex of $F_{(\beta)}$ and therefore
@@
Recall that for all $A\in\mathcal{B}_{\alpha}$, $B\in\mathcal{B}_{\beta}$, we have that $\pi_{1}(A)\subseteq F_{(\alpha)}(v_{\alpha})$, $\pi_{1}(B)\subseteq F_{(\beta)}(v_{\beta})$. Therefore by (12),
@@
If $c=1$, then by the definition of the process $\mathcal{B}_{z}$ is non-empty for each $z\in[k]$. Taking one element in $\mathcal{B}_{z}$ for each $z\in[k]$ yields an independent set (by (13)) of size $k$, which establishes ˜a.
@@
Now suppose $c\geq 2$. For each $z\in[k]$, we fix an independent set $\mathcal{D}_{z}$ of $(F_{i})_{i\in\{2,\ldots,c\}}$-tuples contained in $\mathcal{C}_{z}(v_{z})$ with $|\mathcal{D}_{z}|\geq k^{c-1}$. By ˜13, for each $z\in[k]$ there exists $D_{z}\in\mathcal{D}_{z}$ such that $\{D_{1},\ldots,D_{k}\}$ is independent. For each $z\in[k]$, select some $A_{z}\in\mathcal{B}_{z}$ such that $(\pi_{2}(A_{z}),\ldots,\pi_{c}(A_{z}))=D_{z}$. We now show that $\{A_{1},\ldots,A_{k}\}$ is an independent set of $(F_{i})_{i\in[c]}$-tuples. To this end, let $A:=A_{\alpha}$ and $B:=A_{\beta}$ for some $\alpha,\beta\in[k]$ with $\alpha<\beta$. We must show that $V(\pi_{i}(A))\cap V(\pi_{i}(B))=\emptyset$, for each $i\in[c]$. Since $A\in\mathcal{B}_{\alpha}$ and $B\in\mathcal{B}_{\beta}$, by (13) we have that $V(\pi_{1}(A))\cap V(\pi_{1}(B))=\emptyset$. Since $(\pi_{2}(A),\ldots,\pi_{c}(A))=D_{\alpha}$ and $(\pi_{2}(B),\ldots,\pi_{c}(B))=D_{\beta}$, and since $D_{\alpha}$ and $D_{\beta}$ are independent, we conclude that $V(\pi_{i}(A))\cap V(\pi_{i}(B))=\emptyset$, for each $i\in\{2,\ldots,c\}$. This proves that $\{A_{1},\ldots,A_{k}\}\subseteq\mathcal{A}$ is an independent set, so (a) holds.
Now we consider the case in which $m<k$. First suppose that $c=1$. Note that $|X_{1}|=m\leq k=(\prod_{j=1}^{c-1}2^{j!})\,k^{c!}$ in this case. We claim that $X_{1}:=\{v_{1},\ldots,v_{m}\}$ satisfies ˜a. Suppose the opposite. Then there exists some $A\in\mathcal{A}$ with $X_{1}\cap V(\pi_{1}(A))=\emptyset$. Let $v$ be the root of $\pi_{1}(A)$. The proof splits into two cases depending on whether $\{v_{1},\ldots,v_{m}\}$ contains an ancestor of $v$ in $F_{1}$ or not. If no ancestor of $v$ lies in $\{v_{1},\ldots,v_{m}\}$, then in particular $r_{1}\not\in\{v_{1},\ldots,v_{m}\}$ and $A\in\mathcal{B}_{m}(r_{1})$. This is a contradiction with the fact that the process has terminated. If $\{v_{1},\ldots,v_{m}\}$ contains an ancestor of $v$ in $F_{1}$, fix $z\in[m]$ such that $v_{z}$ is an ancestor of $v$ in $F_{1}$ and is closest to $v$ in $F_{1}$. Consider the moment that the process has chosen $v_{z}$. We have $A\in\mathcal{B}_{z}(u)$, so $v_{z}$ was not a minimal choice for $v_{z}$, contradiction.
Now assume $c\geq 2$. For each $z\in[m]$, let $U_{z}$ be the set of (at most two) children of $v_{z}$ in $F_{(z)}$. Let $u\in U_{z}$. Since $u\in V(F_{(z)})$ we know that $u\not\in\{v_{1},\ldots,v_{z}\}$. Since $u$ was not chosen to be $v_{z}$, we know that $\mathcal{C}_{z}(u)$ has no independent set of size $k^{c-1}$. Let $\mathcal{E}_{z}:=\bigcup_{u\in U_{z}}\mathcal{C}_{z}(u)$. Therefore $\mathcal{E}_{z}$ does not contain an independent set of size $2k^{c-1}$ for each $z\in[m]$.
@@
Next, recall that $F_{1}$ is a tree with root $r_{1}$. If $r_{1}=v_{m}$, then define $\mathcal{E}_{m+1}:=\emptyset$. Otherwise, define $\mathcal{E}_{m+1}:=\mathcal{C}_{m+1}(r_{1})$. (Note that $\mathcal{C}_{m+1}(v)$ is defined for all $v\in V(F_{(z)}(r_{1}))$ so in particular for $v=r_{1}$.) In the latter case, since the process for constructing $v_{1},\ldots,v_{m}$ stopped at iteration $m$, $\mathcal{E}_{m+1}=\mathcal{C}_{m+1}(r_{1})$ does not contain an independent set of size $k^{c-1}$. Therefore $\mathcal{A^{\prime}}:=\bigcup_{z\in[m+1]}\mathcal{E}_{z}$ does not contain an independent set of size $(m+1)2k^{c-1}\leq 2k^{c}$.
@@
Observe that $\mathcal{A}^{\prime}$ is a set of non-trivial $(F_{i})_{i\in\{2,\ldots,c\}}$-tuples. Apply the inductive hypothesis with $c^{\prime}:=c-1$ and $k^{\prime}:=2k^{c}$ to the set $\mathcal{A}^{\prime}$ of non-trivial $(F_{i})_{i\in\{2,\ldots,c\}}$-tuples. Since $\mathcal{A}^{\prime}$ does not contain an independent set of size $k^{\prime}$, this results in outcome ˜a. That is, we obtain a sequence of sets $X_{2},\ldots,X_{c}$ with $X_{i}\subseteq V(F_{i})$ for each $i\in\{2,\ldots,c\}$ such that $\sum_{i=2}^{c}|X_{i}|\leq\ell(2k^{c},c-1)$, and, for each $(A_{2},\ldots,A_{c})\in\mathcal{A}^{\prime}$, there exists some $i\in\{2,\ldots,c\}$ such that $V(A_{i})\cap X_{i}\neq\emptyset$.
Let $X_{1}:=\{v_{1},\ldots,v_{m}\}$. We claim that $X_{1},\ldots,X_{c}$ satisfy the conditions of outcome ˜a.
@@
First we show that, for each $(A_{1},\ldots,A_{c})\in\mathcal{A}$, there exists $i\in[c]$ such that $V(A_{i})\cap X_{i}\neq\emptyset$. If $(A_{2},\ldots,A_{c})\in\mathcal{A}^{\prime}$ then, by the inductive hypothesis, $V(A_{i})\cap X_{i}\neq\emptyset$ for some $i\in\{2,\ldots,c\}$. If $(A_{2},\ldots,A_{c})\notin\mathcal{A}^{\prime}$, then this is because there is no $z\in[m+1]$ such that $(A_{2},\ldots,A_{c})\in\mathcal{E}_{z}$. Let $v$ be the root of $A_{1}$.
First, suppose that $v$ has no $F_{1}$-ancestor in $X_{1}$. Observe that $r_{1}\notin X_{1}$ in this case, and thus $F_{(m+1)}(r_{1})$ and $\mathcal{C}_{m+1}(r_{1})$ are defined. Observe also that $A_{1}\not\subseteq F_{(m+1)}(r_{1})$, which is the component of $F_{(m+1)}=F_{1}-X_{1}$ that contains $r_{1}$, because $(A_{2},\dots,A_{c})\notin\mathcal{E}_{m+1}=\mathcal{C}_{m+1}(r_{1})$. Since $v$ has no $F_{1}$-ancestor in $X_{1}$, $v$ is in $F_{(m+1)}(r_{1})$. Since $A_{1}\not\subseteq F_{(m+1)}(r_{1})$, it must be that $A_{1}$ contains a vertex in $X_{1}$, so $V(A_{1})\cap X_{1}\neq\emptyset$, as required.
[…]
All that remains is to show is that $|X_{i}|\leq(\prod_{j=1}^{c-1}2^{j!})\,k^{c!}$ for each $i\in[c]$. For $i=1$, we have $|X_{1}|=m\leq k$, which satisfies our requirement since $c\geq 1$. For $i\in\{2,\ldots,c\}$, recalling that $k^{\prime}=2k^{c}$, by the inductive hypothesis we obtain
@@
as desired. ∎
Finally, we can complete the proof of Theorem˜3 (restated here for convenience).
@@
By adding a vertex $r$ adjacent to one vertex in each component of $F$, we may assume that $F$ is a tree rooted at $r$. As in the proof of Lemma˜15, we may further assume that $F$ is a binary tree.
Let $\ell:\mathbb{N}^{2}\to\mathbb{N}$ be the function given by Lemma˜15.
The proof is by induction on $c$. When $c=1$, $\mathcal{A}$ is a set of $(F_{i})_{i\in[1]}$-tuples and the result follows, with $\ell^{\star}(k,1)=\ell(k,1)$, from Lemma˜15. Now assume that $c\geq 2$.
Let $A\in\mathcal{A}$ and let $c^{\prime}$ be the number of components of $A$. Thus $c^{\prime}\in[c]$. Define $\Pi(A):=(A_{1},\ldots,A_{c})$ so that $A_{1},\ldots,A_{c^{\prime}}$ are the components of $A$ while $A_{c^{\prime}+1},\ldots,A_{c}$ are the null graphs. (The ordering of the components $A_{1},\ldots,A_{c^{\prime}}$ is chosen arbitrarily.)
For each $\{i,j\}\in\binom{[c]}{2}$ and each $A\in\mathcal{A}$, define $A_{\{i,j\}}$ as the graph obtained from $A$ as follows. If one or both of $\pi_{i}(\Pi(A))$ or $\pi_{j}(\Pi(A))$ is the null graph then $A_{\{i,j\}}:=A$. Otherwise $A_{\{i,j\}}$ is the graph obtained from $A$ by adding the shortest path in $F$ that has one endpoint in $\pi_{i}(\Pi(A))$ and the other endpoint in $\pi_{j}(\Pi(A))$. In either case, $A_{\{i,j\}}$ has at most c1c-1 connected components.
For each $\{i,j\}\in\binom{[c]}{2}$, let $\mathcal{A}_{i,j}:=\{A_{\{i,j\}}:A\in\mathcal{A}\}$. Then $\mathcal{A}_{\{i,j\}}$ is a set of non-null subgraphs of $F$ each having at most c1c-1 components. Now apply the inductive hypothesis to $\mathcal{A}_{\{i,j\}}$. If outcome ˜a holds, then the $k$ pairwise vertex-disjoint members of $\mathcal{A}_{\{i,j\}}$ correspond to $k$ pairwise vertex-disjoint members of $\mathcal{A}$ and there is nothing more to do. Otherwise, outcome ˜b holds and we obtain a set $X_{\{i,j\}}$ with $|X_{\{i,j\}}|\leq\ell^{\star}(k,c-1)$ such that $X_{\{i,j\}}\cap V(A_{\{i,j\}})\neq\emptyset$ for each $A\in\mathcal{A}$.
@@
For each $A\in\mathcal{A}_{0}$, define $\Pi_{0}(A)$ as the $(F_{i})_{i\in[m]}$-tuple in which $\pi_{i}(\Pi_{0}(A))$ is the component of $A$ contained in $F_{i}$ or the null graph if $V(A)\cap V(F_{i})=\emptyset$, for each $i\in[m]$. Since $A$ is non-null, $\Pi_{0}(A)$ is non-trivial. At this point, we can immediately apply Lemma˜15 with $c^{\prime}:=m$ to the set $\{\Pi_{0}(A):A\in\mathcal{A}_{0}\}$ of non-trivial $(F_{i})_{i\in[m]}$-tuples. If outcome ˜a occurs then the resulting set of $k$ independent $(F_{i})_{i\in[m]}$-tuples corresponds to a set of pairwise vertex-disjoint elements of $\mathcal{A}_{0}\subseteq\mathcal{A}$, which satisfies the conditions of ˜a. If outcome ˜b occurs, then we obtain sets $X_{1},\ldots,X_{m}$ with $X_{i}\subseteq V(F_{i})$, for each $i\in[m]$, of total size at most $\ell(k,m)$ such that for each $A\in\mathcal{A}_{0}$, there exists $i\in[m]$ such that $V(A)\cap X_{i}\neq\emptyset$. Then $X:=X_{0}\cup\bigcup_{i\in[m]}X_{i}$ satisfies the conditions of ˜b with
$|X|\leq|X_{0}|+\sum_{i=1}^{m}|X_{i}|\leq\binom{c}{2}\,\ell^{\star}(k,c-1)+\ell(k,m)\leq\binom{c}{2}\,\ell^{\star}(k,c-1)+\ell(k,2\binom{c}{2}\,\ell^{\star}(k,c-1)).$
@@
which is sufficient to establish Theorem˜3. However, with a bit more work, one can prove a bound of the form $\ell^{\star}(k,c)\in O(k^{h(c)})$ for some $h:\mathbb{N}\to\mathbb{N}$, which is polynomial in $k$ for any fixed $c$.
For each $I\in\binom{[m]}{c}$, define the set $\mathcal{A}_{I}:=\{A\in\mathcal{A}_{0}:A\subseteq\bigcup_{i\in I}F_{i}\}$. Since each element in $A$ has at most $c$ components and $V(A)\subseteq V(F_{0})$ for each $A\in\mathcal{A}_{0}$, each such $A$ appears in $\mathcal{A}_{I}$ for at least one $I\in\binom{[m]}{c}$. For each $I\in\binom{[m]}{c}$ and each $A\in\mathcal{A}_{I}$, define $\Pi_{I}(A):=(\pi_{i}(\Pi_{0}(A))_{i\in I}$ as the $(F_{i})_{i\in I}$-tuple obtained by restricting $\Pi_{0}(A)$ to the indices in $I$. For each $I\in\binom{[m]}{c}$, apply Lemma˜15 to the set $\{\Pi_{i}(A):A\in\mathcal{A}_{I}\}$ of $(F_{i})_{i\in I}$-tuples. If outcome ˜a occurs for some $I\in\binom{[m]}{c}$ then the resulting set of $k$ independent $(F_{i})_{i\in I}$-tuples corresponds to a set of pairwise vertex-disjoint elements of $\mathcal{A}_{I}\subseteq\mathcal{A}$, which satisfies the conditions of ˜a. If outcome ˜b occurs for all $I\in\binom{[m]}{c}$, then we obtain a set $X_{I}\subseteq V(F_{0})\subseteq V(F)$ with $|X_{I}|\leq\ell(k,c)$ such that $V(A)\cap X_{I}\neq\emptyset$ for each $A\in\mathcal{A}_{I}$.